OpenStudy (anonymous):

Will Give Medal For Correct Answer Please Show work Subtract: x^2-1/x^2-x-2 - x-1/x-2

OpenStudy (anonymous):

factorize the numerator and denominator on the left ratio

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i don't know how to factor that particular one

OpenStudy (anonymous):

factor out \[x^2-1=?\]

OpenStudy (anonymous):

then factor out\[x^2-x-2=?\]

OpenStudy (anonymous):

yea i don't know how to factor the numerator

OpenStudy (anonymous):

\[x^2-1=(x+1)(x-1)\]

OpenStudy (anonymous):

now, what about the denominator?

OpenStudy (anonymous):

im working on it

OpenStudy (anonymous):

yea nvm i cant figure that one out either, im not realy good at this

OpenStudy (anonymous):

oh wait i got it, is it (x-2)(x+1)?

OpenStudy (anonymous):

good what the the first fraction simplify to?

OpenStudy (anonymous):

x-1/x-2

OpenStudy (anonymous):

good \[{x-1\over x-2}-{x-1\over x-2}=?\]

OpenStudy (anonymous):

0 ?

OpenStudy (anonymous):

@electrokid